Two strings each inclined at an angle of 30 to the horizontal are used to lift a bucket of water weighing 20N. Calculate the tension in each string.
1
Expert's answer
2020-10-27T10:07:11-0400
"2T\cos(\theta/2)=mg"

"T=\frac{mg}{2\cos(\theta/2)}"

"=\frac{20\:\rm N}{2\cos(60^{\circ})}=20\:\rm N."